# Prompt Optimization with EASE? Efficient Ordering-aware Automated Selection of Exemplars

### v1

@@ C.5 Ablation studies for other numbers of $k$
Table 10: Average validation accuracy $\pm$ standard error for $k=50$ and $n=1000$.
@@
AG News Remap 50.0$\pm$2.4 55.0$\pm$2.4 66.7$\pm$1.4
SST5 Reverse 58.3$\pm$1.4 56.7$\pm$1.4 65.0$\pm$2.4
@@
AG News Remap 12.7$\pm$2.6 26.3$\pm$0.7 37.3$\pm$2.4
SST5 Reverse 39.0$\pm$0.5 38.7$\pm$1.0 40.7$\pm$1.0
C.6 Ablation studies for the benefit of having an ordering-aware embedding in EASE
Note that even for the same subset of exemplars, a different ordering leads to different pre-trained embedding from embedding model $h(\cdot)$. In this section, we investigate the impact of capturing the ordering of the exemplars in the exemplar sequence $E$. To contrast the adopted approach, we propose a new embedding that simply averages the embeddings of all exemplars in the chosen subset. So, this embedding will be invariant with regard to the ordering of exemplars and we call it AvgEmb. As demonstrated in Tab. 12 and Tab. 13, adopting an ordering-aware embedding in EASE results in better overall performance as compared to AvgEmd which disregards exemplar ordering.
Table 12: Average validation accuracy $\pm$ standard error for ablation studies of using an embedding without considering order.
@@ C.6 Ablation studies for the benefit of having an ordering-aware embedding in EASE
LR 0% 76.7$\pm$1.4 81.7$\pm$3.6
10% 73.3$\pm$3.6 73.3$\pm$3.6
30% 70.0$\pm$0.0 78.3$\pm$1.4
50% 68.3$\pm$3.6 71.7$\pm$2.7
70% 66.7$\pm$1.4 66.7$\pm$3.6
90% 51.7$\pm$1.4 53.3$\pm$2.7
LP-variant 0% 76.7$\pm$1.4 75.0$\pm$0.0
10% 75.0$\pm$0.0 75.0$\pm$2.4
30% 70.0$\pm$0.0 73.3$\pm$1.4
50% 78.3$\pm$1.4 76.7$\pm$2.7
70% 75.0$\pm$2.4 75.0$\pm$0.0
90% 61.7$\pm$1.4 63.3$\pm$1.4
AG News Remap 0% 53.3$\pm$1.4 53.3$\pm$3.6
10% 50.0$\pm$4.1 56.7$\pm$2.7
30% 51.7$\pm$1.4 51.7$\pm$1.4
50% 50.0$\pm$0.0 56.7$\pm$1.4
70% 58.3$\pm$1.4 51.7$\pm$1.4
90% 53.3$\pm$4.9 55.0$\pm$2.4
SST5 Reverse 0% 46.7$\pm$5.4 50.0$\pm$0.0
10% 50.0$\pm$2.4 50.0$\pm$0.0
30% 45.0$\pm$2.4 41.7$\pm$3.6
50% 48.3$\pm$1.4 43.3$\pm$1.4
70% 41.7$\pm$3.6 45.0$\pm$2.4
90% 31.7$\pm$1.4 31.7$\pm$1.4
# best-performing tasks 11 16
@@
LR 0% 48.7$\pm$1.0 54.0$\pm$4.5
10% 48.7$\pm$2.3 48.3$\pm$1.5
30% 53.0$\pm$4.0 51.0$\pm$1.6
50% 46.0$\pm$4.1 55.0$\pm$1.9
70% 44.3$\pm$5.2 45.7$\pm$1.4
90% 28.7$\pm$2.2 30.0$\pm$3.3
LP-variant 0% 47.7$\pm$1.5 49.3$\pm$2.0
10% 48.0$\pm$1.4 46.0$\pm$0.8
30% 49.3$\pm$1.4 49.3$\pm$1.1
50% 48.0$\pm$1.2 54.3$\pm$2.1
70% 54.3$\pm$1.1 49.0$\pm$3.3
90% 38.3$\pm$0.7 41.3$\pm$1.9
AG News Remap 0% 33.3$\pm$1.9 29.7$\pm$1.9
10% 32.0$\pm$0.5 36.3$\pm$1.0
30% 29.7$\pm$1.0 32.3$\pm$3.7
50% 28.7$\pm$1.2 43.7$\pm$2.8
70% 40.0$\pm$4.1 31.0$\pm$2.5
90% 36.3$\pm$3.3 40.7$\pm$3.1
SST5 Reverse 0% 31.0$\pm$3.7 32.3$\pm$2.8
10% 31.3$\pm$2.4 30.7$\pm$1.9
30% 27.7$\pm$1.2 26.3$\pm$2.2
50% 33.7$\pm$1.2 29.0$\pm$2.9
70% 25.7$\pm$5.9 33.0$\pm$1.4
90% 13.7$\pm$0.3 11.3$\pm$1.2
# best-performing tasks 10 14
C.7 Ablation studies for optimizing exemplars for different black-box LLMs in EASE
@@
Table 14: Average validation accuracy $\pm$ standard error when using our EASE to select exemplars for different target black-box models.
@@
LR 3.3$\pm$1.4 0.0$\pm$0.0 80.0$\pm$2.4
LP-variant 88.3$\pm$1.4 86.7$\pm$1.4 31.7$\pm$1.4
AG News Remap 36.7$\pm$2.7 30.0$\pm$0.0 53.3$\pm$1.4
SST5 Reverse 21.7$\pm$2.7 48.3$\pm$1.4 43.3$\pm$1.4
@@
LR 0.0$\pm$0.0 0.0$\pm$0.0 56.3$\pm$0.7
LP-variant 79.3$\pm$1.4 73.3$\pm$0.7 15.3$\pm$1.7
AG News Remap 20.7$\pm$3.7 26.7$\pm$1.5 37.0$\pm$0.8
SST5 Reverse 12.3$\pm$1.7 30.7$\pm$1.9 19.7$\pm$1.4
C.8 Ablation studies for using different embedding models in EASE
For our main experiments, we use MPNet as the embedding model for our EASE. Here, we use MiniLM [34] and CLIP [26] to obtain the embedding for our EASE, respectively. Tab. 16 and Tab. 17 show that our EASE achieves competitive performance using different embedding models. Therefore, our EASE is general in the sense that different embedding models can be used.
Table 16: Average validation accuracy $\pm$ standard error when using different embedding models for our EASE.
@@ C.8 Ablation studies for using different embedding models in EASE
LR 75.0$\pm$0.0 66.7$\pm$1.4
LP-variant 78.3$\pm$1.4 70.0$\pm$0.0
AG News Remap 46.7$\pm$1.4 45.0$\pm$0.0
SST5 Reverse 43.3$\pm$2.7 46.7$\pm$1.4
@@
LR 49.7$\pm$3.3 45.3$\pm$4.1
LP-variant 48.0$\pm$2.2 51.3$\pm$1.2
AG News Remap 27.0$\pm$2.9 31.0$\pm$1.2
SST5 Reverse 36.7$\pm$1.9 31.7$\pm$1.2
C.9 Ablation studies: Asking GPT to directly select exemplars
@@
Table 18: Average validation accuracy $\pm$ standard error when asking GPT to directly output the best exemplars.
@@
LR 35.0$\pm$6.2 73.3$\pm$3.6
LP-variant 58.3$\pm$2.7 75.0$\pm$2.4
AG News Remap 10.0$\pm$0.0 56.7$\pm$2.7
SST5 Reverse 10.0$\pm$0.0 50.0$\pm$0.0
@@
LR 29.0$\pm$7.4 48.3$\pm$1.5
LP-variant 44.0$\pm$0.8 46.0$\pm$0.8
AG News Remap 12.7$\pm$0.3 36.3$\pm$1.0
SST5 Reverse 10.0$\pm$0.0 30.7$\pm$1.9
C.10 Test accuracies for all tables
We present the test accuracies for all tables presented in the main paper. Note that we report validation accuracies in the main tables because the effectiveness of the optimization strategy is directly reflected by the maximized value of the objective function in (1) at the end of all iterations. Nevertheless, we show all test accuracies here for reference and completeness. The test accuracy can be affected by the difference in the distribution of the validation and test data, which is out of our control. This difference in distribution is significant in our case because the validation set only contains 20 data points (limited by the fact that querying the GPT-3.5 API is expensive). This setup is disadvantageous for EASE because the optimized exemplar is “overfitted” to the validation set. This test performance is expected to improve much if we use a larger validation set (e.g., with 100 validation samples).

### v2

@@ C.5 Ablation studies for other numbers of $k$
Table 10: Average accuracy $\pm$ standard error for $k=50$ and $n=1000$.
@@
AG News Remap 50.0$\pm$2.9 55.0$\pm$2.9 68.3$\pm$1.7
SST5 Reverse 58.3$\pm$1.7 56.7$\pm$1.7 63.3$\pm$1.7
@@
AG News Remap 12.7$\pm$3.2 26.3$\pm$0.9 38.7$\pm$3.2
SST5 Reverse 39.0$\pm$0.6 38.7$\pm$1.2 40.0$\pm$3.5
Another alternative is to allow a range instead of a fixed value of $k$ to be selected. That is, if $k=50$, we can allow the number of exemplars in the prompt to be any integer from 1 to 50. We present the additional results in Tab. 12. Firstly, EASE continues to consistently outperform the strongest baseline Best-of-N. Secondly, we also observe that the prompts with the best performance typically have a large number of exemplars, i.e., close to the max $k$ allowed. Thirdly, using $k=50$ gives better performance than $k=5$. Therefore, including more exemplars in the prompt usually gives a higher performance. However, this comes at the expense of a higher query cost at test time because more tokens are used in the prompt as well.
Table 12: Results for the setting that allows different number exemplars to be selected (i.e., allowing a range instead of a fixed $k$. EASE outperforms Best-of-N. We also observe high-performing prompts usually consist of more exemplars (i.e., larger average $k$).
Tasks $n$ Max $k$ Best-of-N EASE
Acc Avg. best $k$ Acc Avg. best $k$
LR 100 5 60.0$\pm$2.9 5.0$\pm$0.0 76.7$\pm$1.7 4.7$\pm$0.3
LP-variant 100 5 60.0$\pm$0.0 2.7$\pm$0.7 71.7$\pm$1.7 4.3$\pm$0.3
AG News Remap 100 5 31.7$\pm$1.7 3.0$\pm$0.0 46.7$\pm$1.7 5.0$\pm$0.0
SST5 Reverse 100 5 31.7$\pm$1.7 4.3$\pm$0.7 48.3$\pm$1.7 4.7$\pm$0.3
AG News Remap 1000 50 45.0$\pm$2.9 25.0$\pm$0.0 50.0$\pm$2.9 45.0$\pm$1.7
SST5 Reverse 1000 50 60.0$\pm$0.0 38.3$\pm$8.8 61.7$\pm$1.7 36.0$\pm$6.7
C.6 Ablation studies for the query budget
In the main text of the paper, we follow Lin et al. [21] and set the black-box query budget to 165 evaluations. Here, we increase the query budget to 500 iterations to study the effect of the query budget on performance. We show in Tab. 13 that increasing the budget generally improves performance.
Table 13: Average performance when the query budget is increased to 500 iterations. The performance increased as compared to Tab. 2 of the original paper. The gaps between EASE and baselines are still significant.
Tasks (10% noise) Evo Best-of-N EASE
LR 70.0$\pm$0.0 70.0$\pm$0.0 76.7$\pm$3.3
LP-variant 73.3$\pm$1.7 66.7$\pm$1.7 80.0$\pm$0.0
AG News Remap 40.0$\pm$0.0 50.0$\pm$0.0 56.7$\pm$1.7
SST5 Reverse 40.0$\pm$0.0 41.7$\pm$1.7 51.7$\pm$1.7
C.7 Ablation studies for the benefit of having an ordering-aware embedding in EASE
Table 14: Average accuracy $\pm$ standard error for ablation studies of using an embedding without considering order.
@@ C.7 Ablation studies for the benefit of having an ordering-aware embedding in EASE
LR 0% 76.7$\pm$1.7 80.0$\pm$2.9
10% 73.3$\pm$4.4 73.3$\pm$1.7
30% 70.0$\pm$0.0 76.7$\pm$1.7
50% 68.3$\pm$4.4 78.3$\pm$4.4
70% 66.7$\pm$1.7 66.7$\pm$1.7
90% 51.7$\pm$1.7 53.3$\pm$1.7
LP-variant 0% 76.7$\pm$1.7 75.0$\pm$0.0
10% 75.0$\pm$0.0 80.0$\pm$2.9
30% 70.0$\pm$0.0 75.0$\pm$0.0
50% 78.3$\pm$1.7 78.3$\pm$1.7
70% 75.0$\pm$2.9 71.7$\pm$1.7
90% 61.7$\pm$1.7 66.7$\pm$1.7
AG News Remap 0% 53.3$\pm$1.7 50.0$\pm$0.0
10% 50.0$\pm$5.0 51.7$\pm$1.7
30% 51.7$\pm$1.7 55.0$\pm$0.0
50% 50.0$\pm$0.0 55.0$\pm$2.9
70% 58.3$\pm$1.7 58.3$\pm$3.3
90% 53.3$\pm$6.0 53.3$\pm$1.7
SST5 Reverse 0% 46.7$\pm$6.7 50.0$\pm$2.9
10% 50.0$\pm$2.9 48.3$\pm$1.7
30% 45.0$\pm$2.9 46.7$\pm$3.3
50% 48.3$\pm$1.7 46.7$\pm$3.3
70% 41.7$\pm$4.4 45.0$\pm$5.0
90% 31.7$\pm$1.7 31.7$\pm$1.7
# best-performing tasks 9 18
@@
LR 0% 48.7$\pm$1.2 58.0$\pm$2.1
10% 48.7$\pm$2.9 42.3$\pm$0.9
30% 53.0$\pm$4.9 47.7$\pm$5.8
50% 46.0$\pm$5.0 50.7$\pm$2.2
70% 44.3$\pm$6.4 47.0$\pm$3.2
90% 28.7$\pm$2.7 39.0$\pm$7.8
LP-variant 0% 47.7$\pm$1.9 53.0$\pm$2.3
10% 48.0$\pm$1.7 48.0$\pm$2.5
30% 49.3$\pm$1.8 50.7$\pm$1.4
50% 48.0$\pm$1.5 56.7$\pm$0.9
70% 54.3$\pm$1.3 52.0$\pm$1.0
90% 38.3$\pm$0.9 44.3$\pm$1.4
AG News Remap 0% 33.3$\pm$2.3 30.3$\pm$1.2
10% 32.0$\pm$0.6 30.3$\pm$4.4
30% 29.7$\pm$1.2 40.0$\pm$1.0
50% 28.7$\pm$1.4 40.0$\pm$3.1
70% 40.0$\pm$5.0 45.0$\pm$2.0
90% 36.3$\pm$4.1 36.7$\pm$3.2
SST5 Reverse 0% 31.0$\pm$4.6 34.0$\pm$3.0
10% 31.3$\pm$3.0 31.3$\pm$2.6
30% 27.7$\pm$1.4 24.3$\pm$1.8
50% 33.7$\pm$1.4 28.7$\pm$1.4
70% 25.7$\pm$7.2 33.3$\pm$3.8
90% 13.7$\pm$0.3 19.0$\pm$6.1
# best-performing tasks 8 16
Note that even for the same subset of exemplars, a different ordering leads to different pre-trained embedding from embedding model $h(\cdot)$. In this section, we investigate the impact of capturing the ordering of the exemplars in the exemplar sequence $E$. To contrast the adopted approach, we propose a new embedding that simply averages the embeddings of all individual exemplars in the chosen subset. So, this embedding will be invariant with regard to the ordering of exemplars and we call it AvgEmb. As demonstrated in Tab. 14 and Tab. 15, adopting an ordering-aware embedding in EASE results in better overall performance as compared to AvgEmd which disregards exemplar ordering.
Notably, AvgEmd presents a trade-off between efficiency and effectiveness. AvgEmd is advantageous in terms of efficiency as it eliminates the on-the-fly embedding entirely. However, there is no free lunch: Such computational simplification comes at the cost of performance due to the loss of order information. Though not as good, note that AvgEmb still achieves a decent and competitive performance. The practitioners can balance the trade-off and select the most suitable method.
C.8 Ablation studies for optimizing exemplars for different black-box LLMs in EASE
@@
Table 16: Average accuracy $\pm$ standard error when using our EASE to select exemplars for different target black-box models.
@@
LR 1.7$\pm$1.7 3.3$\pm$1.7 83.3$\pm$4.4
LP-variant 90.0$\pm$0.0 90.0$\pm$0.0 31.7$\pm$1.7
AG News Remap 50.0$\pm$0.0 35.0$\pm$2.9 53.3$\pm$4.4
SST5 Reverse 21.7$\pm$1.7 41.7$\pm$1.7 36.7$\pm$1.7
@@
LR 0.0$\pm$0.0 0.0$\pm$0.0 51.3$\pm$3.5
LP-variant 83.7$\pm$1.3 77.7$\pm$1.9 14.3$\pm$1.4
AG News Remap 25.7$\pm$1.9 25.7$\pm$3.0 32.7$\pm$9.4
SST5 Reverse 13.3$\pm$1.4 28.3$\pm$3.8 24.0$\pm$3.0
C.9 Ablation studies for using different embedding models in EASE
For our main experiments, we use MPNet as the embedding model for our EASE. Here, we use MiniLM [43] and CLIP [35] to obtain the embedding for our EASE, respectively. Tab. 18 and Tab. 19 show that our EASE achieves competitive performance using different embedding models. Therefore, our EASE is general in the sense that different embedding models can be used. It might be worth investigating as a future direction to develop or finetune embedding models specifically for the purpose of prompt optimization, such that it captures important aspects of the prompt in the latent space.
Table 18: Average accuracy $\pm$ standard error when using different embedding models for our EASE.
@@ C.9 Ablation studies for using different embedding models in EASE
LR 73.3$\pm$1.7 68.3$\pm$1.7
LP-variant 76.7$\pm$1.7 70.0$\pm$2.9
AG News Remap 43.3$\pm$1.7 46.7$\pm$3.3
SST5 Reverse 43.3$\pm$1.7 45.0$\pm$0.0
@@
LR 51.3$\pm$1.3 50.3$\pm$5.5
LP-variant 56.3$\pm$0.9 49.7$\pm$1.4
AG News Remap 29.3$\pm$2.7 29.3$\pm$0.3
SST5 Reverse 29.7$\pm$6.4 33.3$\pm$4.8
C.10 Ablation studies for the degree of exploration
One important hyperparameter of the NeuralUCB algorithm utilized in the paper is $\nu$ (see (3)), which controls the degree of exploration performed during the optimization process. Our finding indicates that a small $\nu$ in EASE results in better exemplar selection performance. As demonstrated in Tab. 20 and Tab. 21, having a near-zero (i.e., $\nu=0.01$) exploration is the best. This may be attributed to our sub-sampling of the domain space $\Omega$, which is motivated by practically reducing computational complexities, and at the same time inherently serves as a form of exploration. Given the enormous size of the $\Omega$, sub-sampling to $Q_{t}$ likely generates a different exemplar sequences space within each iteration of the algorithm. This observation is further supported by Bastani et al. [2], who found that a greedy algorithm can be rate optimal given sufficient randomness in the observed contexts of the bandits algorithm. Therefore, only a small extent of exploration is required for EASE.
To further support the above observation empirically, we conduct additional experiments with the domain space $Q_{t}$ fixed throughout. When $|Q_{t}|=1000$ and fixed, we observe results in Tab. 22 and Tab. 23 that a larger value of $\nu$ performs better. Thus, exploration is essential when there is no randomness in the context (i.e., the domain space $Q_{t}$). However, given sufficient randomness in our EASE algorithm from sub-sampling, we adopt a small $\nu=0.01$ for optimal performance.
Table 20: Average accuracy $\pm$ standard error for ablation studies of the exploration parameter $\nu$.
Task Noise $\nu=0$ $\nu=0.01$ $\nu=0.1$ $\nu=1$
LP-variant 0% 75.0$\pm$0.0 75.0$\pm$0.0 71.7$\pm$1.7 70.0$\pm$0.0
10% 75.0$\pm$2.9 80.0$\pm$2.9 71.7$\pm$1.7 73.3$\pm$1.7
30% 73.3$\pm$1.7 75.0$\pm$0.0 68.3$\pm$1.7 71.7$\pm$1.7
50% 76.7$\pm$3.3 78.3$\pm$1.7 75.0$\pm$2.9 71.7$\pm$1.7
70% 75.0$\pm$0.0 71.7$\pm$1.7 73.3$\pm$1.7 70.0$\pm$0.0
90% 63.3$\pm$1.7 66.7$\pm$1.7 63.3$\pm$1.7 63.3$\pm$1.7
Table 21: Average test accuracy $\pm$ standard error for ablation studies of the exploration parameter $\nu$.
Task Noise $\nu=0$ $\nu=0.01$ $\nu=0.1$ $\nu=1$
LP-variant 0% 49.3$\pm$2.4 53.0$\pm$2.3 52.0$\pm$1.1 50.0$\pm$3.5
10% 46.0$\pm$1.0 48.0$\pm$2.5 51.3$\pm$2.2 49.3$\pm$1.9
30% 49.3$\pm$1.3 50.7$\pm$1.4 43.7$\pm$3.8 52.3$\pm$2.4
50% 54.3$\pm$2.6 56.7$\pm$0.9 50.3$\pm$2.7 47.0$\pm$1.1
70% 49.0$\pm$4.0 52.0$\pm$1.0 46.0$\pm$2.5 47.0$\pm$1.1
90% 41.3$\pm$2.3 44.3$\pm$1.4 36.7$\pm$2.6 40.3$\pm$1.4
Table 22: Average accuracy $\pm$ standard error for ablation studies of exploration in a fixed domain.
Task Noise $\nu=0$ $\nu=0.01$ $\nu=0.1$ $\nu=1$
LP-variant 0% 60.0$\pm$0.0 66.7$\pm$1.7 68.3$\pm$1.7 66.7$\pm$1.7
10% 60.0$\pm$0.0 61.7$\pm$1.7 66.7$\pm$1.7 65.0$\pm$2.9
30% 58.3$\pm$1.7 58.3$\pm$1.7 65.0$\pm$2.9 65.0$\pm$2.9
50% 60.0$\pm$2.9 65.0$\pm$0.0 61.7$\pm$1.7 66.7$\pm$1.7
70% 50.0$\pm$0.0 51.7$\pm$1.7 58.3$\pm$1.7 61.7$\pm$3.3
90% 36.7$\pm$1.7 36.7$\pm$1.7 45.0$\pm$0.0 48.3$\pm$1.7
Table 23: Average test accuracy $\pm$ standard error for ablation studies of the exploration in a fixed domain.
Task Noise $\nu=0$ $\nu=0.01$ $\nu=0.1$ $\nu=1$
LP-variant 0% 38.3$\pm$0.9 51.0$\pm$3.2 53.7$\pm$3.3 45.0$\pm$1.1
10% 41.7$\pm$2.7 42.7$\pm$1.9 47.0$\pm$1.5 49.0$\pm$2.1
30% 39.3$\pm$2.3 40.3$\pm$3.3 43.3$\pm$1.8 43.0$\pm$1.1
50% 43.3$\pm$2.3 47.7$\pm$0.9 41.7$\pm$2.3 48.7$\pm$2.4
70% 39.7$\pm$1.4 42.0$\pm$3.1 39.7$\pm$1.9 48.7$\pm$0.3
90% 20.7$\pm$0.3 22.3$\pm$0.9 29.0$\pm$3.6 29.0$\pm$1.1
C.11 Ablation studies: Asking GPT to directly select exemplars
@@
Table 24: Average accuracy $\pm$ standard error when asking GPT to directly output the best exemplars.
@@
LR 35.0$\pm$7.6 73.3$\pm$1.7
LP-variant 58.3$\pm$3.3 80.0$\pm$2.9
AG News Remap 10.0$\pm$0.0 51.7$\pm$1.7
SST5 Reverse 10.0$\pm$0.0 48.3$\pm$1.7
@@
LR 29.0$\pm$9.0 42.3$\pm$0.9
LP-variant 44.0$\pm$1.0 48.0$\pm$2.5
AG News Remap 12.7$\pm$0.3 30.3$\pm$4.4
SST5 Reverse 10.0$\pm$0.0 31.3$\pm$2.6
C.12 Additional experiments on tasks with reasoning chains or open-ended generation
Note that the tasks included in the main experiments of the paper already include open-ended generation tasks. Tasks like Auto Categorization, Word Sorting, LR, and LP-variant are beyond simple direct classification-type labels. For example, Auto Categorization requires outputting an open-ended sentence that categorizes the inputs well; LP-variant does open-ended sentence translation following a set of nontrivial rules. Therefore, our EASE demonstrates effectiveness in challenging open-ended generation tasks.
We conduct additional experiments for tasks with reasoning chains, including MATH [15], GSM8K [7], and AQuA-RAT [22]. These tasks are typically harder and involve longer responses. From Tab. 26, EASE works well for these tasks and the advantages are significant.
Table 26: Average performance for tasks involving reasoning chains.
Tasks Evo Best-of-N EASE
MATH [15] 61.7$\pm$1.7 60.0$\pm$0.0 65.0$\pm$0.0
GSM8K [7] 70.0$\pm$2.9 68.3$\pm$1.7 75.0$\pm$0.0
AQuA-RAT [22] 46.7$\pm$1.7 43.3$\pm$1.7 48.3$\pm$1.7
C.13 Additional experiments on more datasets
We conduct experiments on a number of additional real benchmarks. We present results on the benchmarks used in the TEMPERA [49] paper in Tab. 27, including MR, Yelp P., CR, MNLI & MRPC. Note that these tasks are overly simple to distinguish the effectiveness of EASE because all of them achieve above 80% accuracy and are hard to improve further using mere in-context examples.
Therefore, we refer the reviewer to more complicated benchmarks in the main paper. Additionally, real tasks in Sec. C.12 also demonstrated good performance of EASE on MATH, GSM8K & AQuA-RAT with Chain-of-Thought reasoning, making the baseline comparisons in our paper more convincing.
Table 27: Average performance on benchmarks used in TEMPERA [49].
Tasks Evo Best-of-N EASE
MR 95.0$\pm$0.0 95.0$\pm$0.0 95.0$\pm$0.0
Yelp P. 95.0$\pm$0.0 95.0$\pm$0.0 95.0$\pm$0.0
CR 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
MNLI 81.7$\pm$1.7 80.0$\pm$0.0 81.7$\pm$1.7
MRPC 81.7$\pm$1.7 81.7$\pm$1.7 81.7$\pm$1.7
C.14 Open-weight models for reproducibility
For reproducibility’s sake, we conduct additional experiments with open-weight models. We use a representative (Meta’s newest) model Llama-3.1-8B-Instruct as the target model. We present the results in Tab. 28. Importantly, the results are consistent with the original conclusions we drew for the black-box models in the paper. The model is also much smaller than GPT-3.5, hence making it easier for others to deploy and reproduce the results.
Table 28: Average performance on Meta’s newest open-weight model Llama-3.1-8B-Instruct.
Tasks (with 50% noise) Evo Best-of-N EASE
LR 11.7$\pm$1.7 15.0$\pm$2.9 35.0$\pm$2.9
LP-variant 11.7$\pm$1.7 15.0$\pm$0.0 25.0$\pm$0.0
AG News Remap 31.7$\pm$1.7 31.7$\pm$1.7 33.3$\pm$1.7
SST5 Reverse 31.7$\pm$3.3 30.0$\pm$0.0 28.3$\pm$3.3
C.15 Test accuracies for all tables
In Tab. 29, Tab. 30, Tab. 31 and Tab. 32, we present the test accuracies for all tables presented in the main paper. Note that we report validation accuracies in the main tables because the effectiveness of the optimization strategy is directly reflected by the maximized value of the objective function in (1) at the end of all iterations. Nevertheless, we show all test accuracies here for reference and completeness. The test accuracy can be affected by the difference in the distribution of the validation and test data, which is out of our control. This difference in distribution is significant in our case because the validation set only contains 20 data points (limited by the fact that querying the GPT-3.5 API is expensive). This setup is disadvantageous for EASE because the optimized exemplar is “overfitted” to the validation set. This test performance is expected to improve much if we use a larger validation set (e.g., with 100 validation samples).